OpenStudy (anonymous):

help please...prove sin x/sec squared x-1=cosx cotx

OpenStudy (luigi0210):

Do you know your identities?

OpenStudy (anonymous):

yes but im a little confuse how to begin

OpenStudy (luigi0210):

Can you clarify your equation, first? Use the Equation editor.

OpenStudy (anonymous):

\[sinx/\sec ^{2}x-1=cosx cotx\]

OpenStudy (anonymous):

Is that okay?

OpenStudy (luigi0210):

Is the 1 with sec^2x or is it separate?

OpenStudy (luigi0210):

Is it: \[\LARGE \frac{sinx}{sec^2x}-1=cosxcotx\] Or: \[\LARGE \frac{sinx}{sec^2x-1}=cosxcotx\]

OpenStudy (anonymous):

the second one

OpenStudy (luigi0210):

We can use the identity: \[\LARGE tan^2(t) + 1 = sec^2(t)\]

OpenStudy (luigi0210):

So if we subtract the 1, we get: \[\LARGE tan^2x=sec^2x-1\] And we can sub that in instead

OpenStudy (anonymous):

I got it thank you

OpenStudy (luigi0210):

Yup :)